(5-y)+(6-y)-(5-y)=0 what does y equal???
Found 2 solutions by MathGrace, hyh_124:
Answer by MathGrace(6)

First you have to distribute any minus or pluses you have. You can re-write the equation as 5-y+6-y-5+y=0. Then you should combine like terms on the left and right. 5+6-5=6 and -y-y+y=-y. After you've combined like terms you now have 6-y=0. Subtract a 6 from both sides to get -y=-6, then multiply both sides by -1 to make y positive. (-1)(-y)=(-1)(-6) => y=6

Answer by hyh_124(6)

first solve the brackets (BODMAS)
so will have 5-y+6-y-5+y=0
=>5+6-5-y+y-y=0
=>11-5-2y+y=0
=>6-y=0
=> -y=-6 (cross out the negative signs)
=> y=6
Check

(5-y)+(6-y)-(5-y)=6
(5-6)+(6-6)-(5-6)=6
5-6+6-6-5+6=6
5-6-5=6
6=6